OpenStudy (anonymous):

Matt has an extensive CD collection. One-fifth of his CDs are rock music and 3/10 of his CDs are classical music. What fraction of his CDs are types of music other than rock or classical?

OpenStudy (nurali):

solve the fraction \[\frac{ 1 }{ 1 }-\frac{ 1 }{ 5 }-\frac{ 3 }{ 10}\]

OpenStudy (anonymous):

-3/-14 ??

OpenStudy (nurali):

No it is 5/10 is equal to 1/2.